When a 3.08 g sample of liquid octane (C8H18) is burned in a bomb calorimeter, the temperature of the calorimeter rises by 26.9 oC. The heat capacity of the calorimeter, measured in a separate experiment, is 6.22 kJ/∘C . The calorimeter also contains 3.00 kg of water, specific heat capacity of 4.18 J/g°C. Determine the heat of combustion of octane in units of kJ/mol octane.
Answer #1

mass of octane = 3.08 g

moles of octane = 3.08 / 114.2285 = 0.02696 mol

calorimeter heat capacity = 6.22 kJ / oC

temperature rise = 26.9 oC

Q = m Cp dT + Cp dT

   = 3.00 x 4.18 x 26.9 + 6.22 x 26.9

= 504.6 kJ

delta H = Q / n

             = 504.6 / 0.02696

delta H = -1.87 x 10^4 kJ/mol